OpenStudy (anonymous):

WILL MEDAL PLEASE HELP How many solutions are there to the following system of equations? 3x – 9y = 0 –x + 3y = –3 A 0 B 1 C 2 D infinitely many

OpenStudy (youngsuperman):

c

OpenStudy (mathmate):

Example for a system with no solution: 5x+3y=4 -5x-3y=-6 Example for a system with a infinite number of solutions: 5x+3y=4 -5x-3y=-4

OpenStudy (anonymous):

Thank you @mathmate I used a graphing calculator to help, since I'm timed, but it stilll confused me until now, THANKS BUNCHES!!!

ganeshie8 (ganeshie8):

graphing is a good idea too so what did you notice on graph ?

OpenStudy (anonymous):

I noticed the lines didn't cross each other which means that there are no solutions. @ganeshie8

ganeshie8 (ganeshie8):

Perfect !

OpenStudy (anonymous):

Thank you! @ganeshie8 :D